If a firm's demand function is p=100-q,what quantities need to be sold to bring in a total revenue of

1
Expert's answer
2021-11-01T19:23:08-0400

Complete Question

A firm faces the demand schedule q = 400 - 2p - p². What price does it need to charge to sell 100 units?



If a firm's demand function is P = 100 - q, what quantities need to be sold to bring in a total revenue of £100



Solution:

Q =400 – 2p – p2

Q = 100 units

100 = 400 – 2p – p2

P = 16.35

The firm needs to charge a price of 16.35 to sell 100 units.

Demand function: P = 100 – q

TR = P "\times" Q

100 = (100 – q) q

100 = 100q – q2

100q – q2 – 100 = 0

-q2 + 100q – 100 = 0

Solve through quadratic equation:

Q = 98.99 units


The quantities needed to be sold to bring in total revenue of £100 = 98.99 units